Suppose the a firm has the production Q=2K^0.5L^0.5
Assume that the capital is fixed at nine units (ie K=9?). If the priceof output p is 6 per unit and wage rate w is Rs 2 per unit, determined the optimal or maximizing rate of labour to be hired. What labour rate is optimal if the wage rate increased to Rs 3 per unit
1
Expert's answer
2019-05-20T16:16:07-0400

If the price of output p is 6 per unit and wage rate w is Rs 2 per unit, then the optimal or maximizing rate of labour to be hired.is:

K/L = r/w,

9/L = r/2,

L = 18/r, where r is the price of capital.

If the wage rate increased to Rs 3 per unit, then:

K/L = r/w,

9/L = r/3,

L = 27/r.